# FedLPPA: Learning Personalized Prompt and Aggregation for Federated Weakly-supervised Medical Image Segmentation

### v1

@@ Abstract
Federated learning (FL) effectively mitigates the data silo challenge brought about by policies and privacy concerns, implicitly harnessing more data for deep model training. However, traditional centralized FL models grapple with diverse multi-center data, especially in the face of significant data heterogeneity, notably in medical contexts. In the realm of medical image segmentation, the growing imperative to curtail annotation costs has amplified the importance of weakly-supervised techniques which utilize sparse annotations such as points, scribbles, etc. A pragmatic FL paradigm shall accommodate diverse annotation formats across different sites, which research topic remains under-investigated. In such context, we propose a novel personalized FL framework with learnable prompt and aggregation (FedLPPA) to uniformly leverage heterogeneous weak supervision for medical image segmentation. In FedLPPA, a learnable universal knowledge prompt is maintained, complemented by multiple learnable personalized data distribution prompts and prompts representing the supervision sparsity. Integrated with sample features through a dual-attention mechanism, those prompts empower each local task decoder to adeptly adjust to both the local distribution and the supervision form. Concurrently, a dual-decoder strategy, predicated on prompt similarity, is introduced for enhancing the generation of pseudo-labels in weakly-supervised learning, alleviating overfitting and noise accumulation inherent to local data, while an adaptable aggregation method is employed to customize the task decoder on a parameter-wise basis. Extensive experiments on three distinct medical image segmentation tasks involving different modalities underscore the superiority of FedLPPA, with its efficacy closely parallels that of fully supervised centralized training. Our code and data will be available at https://github.com/llmir/FedLPPA.
@@ I Introduction
Concurrently, the pressing need to reduce annotation costs has heightened the focus on weakly supervised segmentation (WSS). WSS techniques typically leverage sparse granularity supervision (like points, boxes, scribbles, or blocks, as shown in Fig. 1B) through novel loss function designing, consistency learning, adversarial learning, or data synthesis [21, 22, 23, 24]. Integrating WSS into FL can further reduce the annotation cost per site while leveraging the benefits of FL. However, there is little research on WSS in pFL. A more challenging but practical setup is to allow for different sites to possess different forms of weak labels, which further introduces label heterogeneity. In such context, it is highly desirable to design a unified and easy-to-deploy pFL framework for WSS settings.
@@
  • Our FedLPPA establishes consistent superiority over general FL and state-of-the-art pFL methods in extensive experiments on three different medical image segmentation tasks involving different modalities.
@@ II-A Federated Learning for Medical Image Analysis
II-B Weakly-supervised Segmentation
@@ III Method
In this section, we first briefly overview our FedLPPA framework, then detail the Tri-prompt Dual-attention Fusion (TDF) module and the Prompt similarity Dual-decoder with Learnable Aggregation (PDLA) mechanism, which are the two key components of FedLPPA. The former leverages in-context learning to enable the task decoders and heads to perceive common knowledge across multiple centers, local data distribution, and supervision granularity for better adaptive adjustments. The latter personalizes the auxiliary decoder’s parameters from sites with similar local distributions to mitigate noise accumulation and generate superior pseudo-labels, while learnable aggregation further personalizes the parameters of the dual-decoder. Additionally, we describe preprocessing operations for bounding box data to align with other supervision forms. The overall framework of FedLPPA is provided in Fig. 2.
Figure 2: Schematic illustration of the proposed FedLPPA framework and its key components. Note that the segmentation heads and the MLP block are incorporated within the decoders and are not depicted independently, except in the upper left panel of the scheme.
@@ III-A FedLPPA Paradigm Overview
With the support of a central server, FedLPPA is designed to establish a personalized model denoted as $\{{\theta}_{i}\}^{N}_{i=1}$ for each client/site without the necessity of data sharing, as opposed to constructing a single shared model in gFL methods. These models not only incorporate beneficial information from other sites but also adaptively fit to the corresponding local data distribution and supervision form. Assume there are $N$ sites with private data $\{\mathcal{D}_{1},...,\mathcal{D}_{N}\}$ where $\mathcal{D}_{i}=\{(\mathbf{X}_{i},\mathbf{Y}_{i})\}_{i=1}^{N}$ with $\mathbf{X}_{i}$ and $\mathbf{Y}_{i}$ respectively denoting the data and labels of the $i^{th}$ site. The data distributions vary, i.e., $\mathcal{P}(\mathcal{D}_{i})\neq\mathcal{P}(\mathcal{D}_{j})$, and the annotation forms of $\{\mathbf{Y}_{i}\}_{i=1}^{N}$ may differ, including types such as point, scribble, bounding box and block. In FedLPPA, the global model and local models share the same network architecture, which is a U-shape network with an additional auxiliary decoder and a TDF module inserted in the middle. Furthermore, a multilayer perceptron (MLP) block links the TDF module with the main decoder’s segmentation head (as shown in Fig. 2A), enabling more direct transfer of context-injected features and enhancing flexibility in parameter optimization. We set the encoder and the TDF module as the globally shared parts, denoted by parameter $\vartheta$. The main decoder (with MLP, parameter $\phi$) and the auxiliary decoder (parameter $\bar{\phi}$) are configured as personalized components. The training process of FedLPPA involves $T$ rounds of alternating updates between clients and server. In the $t^{th}$ round, all clients receive the same aggregated global parameters ($\vartheta_{G}^{t-1}$ and $\phi_{G}^{t-1}$) from the server, as well as customized auxiliary decoder parameter ($\bar{\phi}_{G,i}^{t-1}$) obtained according to the selected strategy. The global part of the local model is initialized using $\vartheta_{G}^{t-1}$, while the personalized parts are initialized with $\hat{\phi}_{i}^{t}$ and $\hat{\bar{\phi}}_{i}^{t}$ which are derived using the learnable aggregation (LA) mechanism respectively based on (${\phi}_{G}^{t-1}$,${\phi}_{i}^{t-1}$) and ($\bar{\phi}_{G,i}^{t-1}$,$\bar{\phi}_{i}^{t-1}$). Each client updates its local model via optimizing the local objective with local data $\mathcal{D}_{i}$ and the manually set annotation sparsity prompt (ASP) $p_{S,i}$ used in TDF,
@@
where $\omega^{i}=|\mathcal{D}_{i}|/\sum_{j=1}^{N}|\mathcal{D}_{j}|$ is the aggregation weight and $|\mathcal{D}_{i}|$ is the sample size of the $i^{th}$ site. The server-side auxiliary decoder parameters $\bar{\phi}_{G,i}^{t}$ will be introduced in Sec. III-C.
@@ III-B Tri-prompt Dual-attention Fusion (TDF) Module
In-context learning is designed to empower networks to adaptively optimize for various contexts by employing conditions or prompts that encapsulate external factors such as scene, task, or modality. Unlike the resource-intensive visual prompts or the rigid encodings of pretrained CLIP outputs or one-hot codes [43, 20, 16], which either burden computation and privacy or miss cross-client insights, we employ a triple-latent-prompt strategy. This approach integrates universal knowledge, data distribution, and annotation sparsity, effectively merging learnable parameters with predefined one-hot encodings for a more nuanced and efficient response. Concretely, all clients jointly maintain two learnable parameters: the universal knowledge prompt (UKP) $p_{U}\in\mathbb{R}^{1\times H\times W}$ and the data distribution prompt (DDP) $p_{D}\in\mathbb{R}^{N\times H\times W}$, with $N$ denoting the number of clients, and $H$ and $W$ respectively representing the height and width of the encoder’s output feature $F$. The federation also provides $n$ one-hot annotation sparsity prompts (ASP) $p_{S}$, specifically set to 3 in this work, to discern between point, scribble, and block levels of sparsity, with box labels preprocessed into one of these three categories.
As illustrated in Fig. 2B, for the $i^{th}$ client, the local prompt $p_{i}$ is constructed by concatenating the shared UKP $p_{U}$, the locally selected DDP $p_{D,i}\in\mathbb{R}^{1\times H\times W}$ from the channel dimension, and the dimensionally expanded ASP $p_{S,i}\in\mathbb{N}^{3\times H\times W}$, denoted as $p_{i}=\mathrm{cat}(p_{U},p_{D,i},p_{S,i})$. Given an encoded feature $F$, it is concatenated with $p_{i}$ and then passed through two convolutional blocks to obtain a preliminary fused feature $\hat{F}$. To further fuse and enhance features, we follow the typical dual-attention mechanism [44] in both spatial and channel dimensions. With $\otimes$ denoting the matrix multiplication operation, the spatial attention value $s_{ji}$ measuring the impact of the $i^{th}$ location on the $j^{th}$ location is calculated as
@@
$\tilde{F}=\gamma_{s}\cdot F_{s}+\hat{F}+\gamma_{c}\cdot F_{c}+\hat{F},$ (5)
@@ III-C PDLA Mechanism and WSS Objective
In WSS, an effective strategy is to generate reliable pseudo-proposals for unlabeled areas. To reduce local computational overhead, we forgo the multi-level tree affinity approach [16]. An alternative is self-training using the model’s own predictions, which is nevertheless prone to update stagnation and noise accumulation [38, 39]. To mitigate these issues, we adopt a dual-decoder strategy wherein an auxiliary branch incorporates beneficial information from other centers. In the PDLA mechanism, parameter acquisition for both decoders involves two stages: aggregation or selection at the server side, and learnable aggregation at the local end.
@@
To stabilize the training of the dual-decoder and facilitate fine-grained personalized knowledge transfer, we employ a learnable aggregation (LA) approach akin to residual learning to learn an aggregation weight matrix (for each decoder at the client end), which is used to personalize decoder parameters on an element-wise basis. LA is formulated as
@@
where $W_{M,i}$ and $W_{A,i}$ are learnable weight matrices, and $\sigma(w)=\max(0,\min(1,w)),\forall w\in\{W_{M,i},W_{A,i}\}$ is utilized to ensure the weight matrices’ elements are within the range of $[0,1]$. $W_{M,i}$ and $W_{A,i}$ are initialized as identity matrices, and then iteratively get updated during training on local data
@@
where $\eta$ is the learning rate. During the update process of $W_{M,i}$ and $W_{A,i}$, Eqs. (10-11) and (12-13) are alternately executed for the updates of $(\hat{\phi}_{i}^{t},\hat{\bar{\phi}}_{i}^{t})$ and $(W_{M,i},W_{A,i})$. Upon convergence, the initial parameters $(\hat{\phi}_{i}^{t},\hat{\bar{\phi}}_{i}^{t})$ for the dual-decoder in each federated round are obtained, followed by the training of the local model using Eq. (1). Note that when $t>2$, only a few iterations are conducted to acquire $(W_{M,i},W_{A,i})$.

### v2

@@ Abstract
Federated learning (FL) effectively mitigates the data silo challenge brought about by policies and privacy concerns, implicitly harnessing more data for deep model training. However, traditional centralized FL models grapple with diverse multi-center data, especially in the face of significant data heterogeneity, notably in medical contexts. In the realm of medical image segmentation, the growing imperative to curtail annotation costs has amplified the importance of weakly-supervised techniques which utilize sparse annotations such as points, scribbles, etc. A pragmatic FL paradigm shall accommodate diverse annotation formats across different sites, which research topic remains under-investigated. In such context, we propose a novel personalized FL framework with learnable prompt and aggregation (FedLPPA) to uniformly leverage heterogeneous weak supervision for medical image segmentation. In FedLPPA, a learnable universal knowledge prompt is maintained, complemented by multiple learnable personalized data distribution prompts and prompts representing the supervision sparsity. Integrated with sample features through a dual-attention mechanism, those prompts empower each local task decoder to adeptly adjust to both the local distribution and the supervision form. Concurrently, a dual-decoder strategy, predicated on prompt similarity, is introduced for enhancing the generation of pseudo-labels in weakly-supervised learning, alleviating overfitting and noise accumulation inherent to local data, while an adaptable aggregation method is employed to customize the task decoder on a parameter-wise basis. Extensive experiments on four distinct medical image segmentation tasks involving different modalities underscore the superiority of FedLPPA, with its efficacy closely parallels that of fully supervised centralized training. Our code and data will be available at https://github.com/llmir/FedLPPA.
@@ I Introduction
Concurrently, the pressing need to reduce annotation costs has heightened the focus on weakly supervised segmentation (WSS). WSS techniques typically leverage sparse granularity supervision (like points, boxes, scribbles, or blocks, as shown in Fig. 1B) through novel loss function designing, consistency learning, adversarial learning, or data synthesis [21, 22, 23, 24]. Integrating WSS into FL can further reduce the annotation cost per site while leveraging the benefits of FL. However, there is little research on WSS in pFL. A more challenging but practical setup is to allow for different sites to possess different forms of weak labels, which further introduces label/supervision heterogeneity. In this setting, the challenges include data heterogeneity, label heterogeneity, and insufficient supervision signals. These issues collectively lead to uncoordinated model training rates or progress across multiple centers in FL, complicating interaction and coordination. In such context, it is highly desirable to design a unified and easy-to-deploy pFL framework for WSS settings.
@@
  • Our FedLPPA establishes consistent superiority over general FL and state-of-the-art pFL methods in extensive experiments on four different medical image segmentation tasks involving different modalities.
This paper extends our FedICRA [16] in several aspects:
  • Discarding the single one-hot prompt, which fails to finely cope with diverse training conditions and misses cross-client insights, we introduce a design using triple latent prompts, two of which are learnable. This allows the model and prompts to more finely perceive and coordinate various training conditions/contexts, while maintaining and representing inter-client common knowledge and relationship. Additionally, we perform multiple fusion of prompts and image features to better enhance the features.
  • Forgoing the multi-level tree affinity WSS paradigm, we propose a PDLA paradigm, which reduces computational overhead. By leveraging prompts to personalize the selection or aggregation of auxiliary decoders, this approach effectively gathers useful information from similar clients and reduces interference. This strategy assists in generating higher-quality and more diverse pseudo-labels, thereby enhancing the segmentation performance.
  • We validate the performance of FedLPPA in a broader range of medical image segmentation scenarios, including endoscopy and MRI. Additionally, we conduct more ablation and analysis experiments to further scrutinize the efficacy of the proposed framework.
@@ II-A Federated Learning for Medical Image Analysis
II-B Singular/Hybrid Weakly-supervised Segmentation
@@ III Method
Figure 2: Schematic illustration of the proposed FedLPPA framework. Note that the segmentation heads and the MLP block are incorporated within the decoders and are not depicted independently, except in the left panel of the scheme. The auxiliary decoder aggregation on the right panel builds its basis on the selection of the Prompt Similarity Aggregated strategy.
In this section, we first briefly overview our FedLPPA framework, then detail the Tri-prompt Dual-attention Fusion (TDF) module and the Prompt similarity Dual-decoder with Learnable Aggregation (PDLA) mechanism, which are the two key components of FedLPPA. The former leverages in-context learning to enable the task decoders and heads to perceive common knowledge across multiple centers, local data distribution, and supervision granularity for better adaptive adjustments. The latter personalizes the auxiliary decoder’s parameters from sites with similar local distributions to mitigate noise accumulation and generate superior pseudo-labels, while learnable aggregation further personalizes the parameters of the dual-decoder. Additionally, we describe preprocessing operations for bounding box data to align with other supervision forms. The overall framework and key components of FedLPPA are respectively provided in Fig. 2 and Fig. 3.
@@ III-A FedLPPA Paradigm Overview
With the support of a central server, FedLPPA is designed to establish a personalized model denoted as $\{{\theta}_{i}\}^{N}_{i=1}$ for each client/site without the necessity of data sharing, as opposed to constructing a single shared model in gFL methods. These models not only incorporate beneficial information from other sites but also adaptively fit to the corresponding local data distribution and supervision form. Assume there are $N$ sites with private data $\{\mathcal{D}_{1},...,\mathcal{D}_{N}\}$ where $\mathcal{D}_{i}=\{(\mathbf{X}_{i},\mathbf{Y}_{i})\}_{i=1}^{N}$ with $\mathbf{X}_{i}$ and $\mathbf{Y}_{i}$ respectively denoting the data and labels of the $i^{th}$ site. The data distributions vary, i.e., $\mathcal{P}(\mathcal{D}_{i})\neq\mathcal{P}(\mathcal{D}_{j})$, and the annotation forms of $\{\mathbf{Y}_{i}\}_{i=1}^{N}$ may differ, including types such as point, scribble, bounding box and block. In FedLPPA, the global model and local models share the same network architecture, which is a U-shape network with an additional auxiliary decoder and a TDF module inserted in the middle. Furthermore, a multilayer perceptron (MLP) block links the TDF module with the main decoder’s segmentation head (as shown in Fig. 2A), enabling more direct transfer of context-injected features and enhancing flexibility in parameter optimization. We set the encoder and the TDF module (including the learnable prompts) as the globally shared parts, denoted by parameter $\vartheta$. The main decoder (with MLP, parameter $\phi$) and the auxiliary decoder (parameter $\bar{\phi}$) are configured as personalized components. The training process of FedLPPA involves $T$ rounds of alternating updates between clients and server. In the $t^{th}$ round, all clients receive the same aggregated global parameters ($\vartheta_{G}^{t-1}$ and $\phi_{G}^{t-1}$) from the server, as well as customized auxiliary decoder parameter ($\bar{\phi}_{G,i}^{t-1}$) obtained according to the selected strategy. The global part of the local model is initialized using $\vartheta_{G}^{t-1}$, while the personalized components are initialized with $\hat{\phi}_{i}^{t}$ and $\hat{\bar{\phi}}_{i}^{t}$, respectively. These parameters ($\hat{\phi}_{i}^{t}$ and $\hat{\bar{\phi}}_{i}^{t}$) are derived using the Learnable Aggregation (LA) mechanism based on the parameters distributed by the server and the local parameters from the previous round (i.e., (${\phi}_{G}^{t-1}$,${\phi}_{i}^{t-1}$) and ($\bar{\phi}_{G,i}^{t-1}$,$\bar{\phi}_{i}^{t-1}$), respectively). Each client updates its local model via optimizing the local objective with local data $\mathcal{D}_{i}$ and the manually set annotation sparsity prompt (ASP) $p_{S,i}$ used in TDF,
@@
where $\omega^{i}=|\mathcal{D}_{i}|/\sum_{j=1}^{N}|\mathcal{D}_{j}|$ is the aggregation weight and $|\mathcal{D}_{i}|$ is the sample size of the $i^{th}$ site. The server-side auxiliary decoder parameters $\bar{\phi}_{G,i}^{t}$ will be introduced in Sec. III-C. Algorithm 1 provides a detailed description of FedLPPA’s training process.
Input: $\{\mathcal{D}_{i}\}_{i=1}^{N}$: training data of $N$ clients; $T$: total federation rounds; $p_{S,i}$: ASP
Output: $\{\theta_{i}\}_{i=1}^{N}$: personalized models of $N$ clients
1
[…]
3 Server sends $\theta_{G}^{0}$ to all clients, $\theta_{i}^{0}\leftarrow\theta_{G}^{0}$
4 Client $i$ selects $p_{D,i}^{0}$, $p_{S,i}$ and initializes $W_{M,i}$, $W_{A,i}$
5 if $\mathcal{D}_{i}$ uses box label then
6       Perform predefined label preprocessing
7
8 end if
9for $t=1:T$ do
10       while $W_{M,i}$, $W_{A,i}$ do not converge do
11             Train $W_{M,i}$, $W_{A,i}$ with Eqs. (12-13) $\triangleright\mathbf{LA}$
12             Clip $W_{M,i}$, $W_{A,i}$ to $[0,1]$
13
14       end while
15      Obtain parameters $\hat{\phi}_{i}^{t}$, $\hat{\bar{\phi}}_{i}^{t}$ with Eqs. (10-11)
[…]
18       Server obtains $\vartheta_{G}^{t}$, $\phi_{G}^{t}$ with Eq. (2)
19       Calculate affinity matrix $A$ with Eqs. (7-8)
[…]
21       Distribute $\{\bar{\phi}_{G,i}^{t}\}_{i=1}^{N}$ to corresponding clients
22
23 end for
24return $\{\theta_{i}\}_{i=1}^{N}$
Algorithm 1 FedLPPA
Figure 3: Detailed illustration of the Tri-prompt Dual-attention Fusion (TDF) module and the Learnable Aggregation (LA) for dual-decoder mechanism. The term ‘context-injected’ denotes that the training conditions or contextual information have been integrated into the original image features, thereby resulting in enhanced features.
@@ III-B Tri-prompt Dual-attention Fusion (TDF) Module
In-context learning is designed to empower networks to adaptively optimize for various contexts by employing conditions or prompts that encapsulate external factors such as scene, task, or modality. Unlike the resource-intensive visual prompts or the rigid encodings of pretrained CLIP outputs or one-hot codes [43, 20, 16], which either burden computation and privacy or miss cross-client insights, we employ a triple-latent-prompt strategy. This approach integrates universal knowledge, data distribution, and annotation sparsity, effectively merging learnable parameters with predefined one-hot encodings for a more nuanced and efficient response. Concretely, all clients jointly maintain two learnable parameters: the universal knowledge prompt (UKP) $p_{U}\in\mathbb{R}^{1\times H\times W}$ and the data distribution prompt (DDP) $p_{D}\in\mathbb{R}^{N\times H\times W}$, with $N$ denoting the number of clients, and $H$ and $W$ respectively representing the height and width of the encoder’s output feature $F$. The federation also provides $n$ one-hot annotation sparsity prompts (ASP) $p_{S}$, specifically set to 3 in this work, to discern between sparse (e.g., point), medium (e.g., scribble), and dense (e.g., block) levels of sparsity, with box labels preprocessed into one of these three categories. ASP is designed to function as a switch to assist the model in perceiving the degree of annotation/supervision sparsity. It facilitates the coordination of training among multi-center models that are trained with supervision signals of varying sparsity levels.
As illustrated in Fig. 2B, for the $i^{th}$ client, the local prompt $p_{i}$ is constructed by concatenating the shared UKP $p_{U}$, the locally selected DDP $p_{D,i}\in\mathbb{R}^{1\times H\times W}$ from the channel dimension, and the dimensionally expanded ASP $p_{S,i}\in\mathbb{N}^{3\times H\times W}$, denoted as $p_{i}=\mathrm{cat}(p_{U},p_{D,i},p_{S,i})$. As suggested by previous works [19, 27, 44], directly adding or concatenating features with prompts may result in suboptimal performance. Rather, employing designs such as convolutional layers or some forms of attention mechanisms to effectively fuse and enhance both elements is more favorable. Given an encoded feature $F$, it is concatenated with $p_{i}$ and then passed through two convolutional blocks to obtain a preliminary fused feature $\hat{F}$. To further fuse and enhance features, we follow the typical dual-attention mechanism [45] in both spatial and channel dimensions. With $\otimes$ denoting the matrix multiplication operation, the spatial attention value $s_{ji}$ measuring the impact of the $i^{th}$ location on the $j^{th}$ location is calculated as
@@
${\tilde{F}=(\gamma_{s}\cdot F_{s}+\hat{F})+(\gamma_{c}\cdot F_{c}+\hat{F}),}$ (5)
@@ III-C PDLA Mechanism and WSS Objective
In WSS, an effective strategy is to generate reliable pseudo-proposals for unlabeled areas. To reduce local computational overhead, we forgo the multi-level tree affinity approach [16]. An alternative approach involves self-training using the model’s own predictions. However, since weak labels typically lack direct supervision at boundary areas, models are prone to being constrained by their own predictions, especially in uncertain regions, which may lead to update stagnation or noise accumulation [38, 39]. To mitigate these issues, we adopt a dual-decoder strategy wherein an auxiliary branch incorporates beneficial information from other centers. This decoder provides predictions from a different or slightly different perspective. Predictions from the main decoder and the auxiliary decoder are then ensembled to yield pseudo labels of higher quality and more diversity. In the PDLA mechanism, parameter acquisition for both decoders involves two stages: aggregation or selection at the server side, and learnable aggregation at the local end.
@@
To stabilize the training of the dual-decoder, to avoid unsuitable personalization due to coarse-grained aggregation, and to facilitate fine-grained and personalized knowledge transfer, we employ a Learnable Aggregation (LA) approach akin to residual learning. This approach involves learning an aggregation weight matrix for each decoder at the client end, which is then used to personalize the decoder parameters on an element/parameter-wise basis. LA is formulated as
@@
where $W_{M,i}$ and $W_{A,i}$ are learnable weight matrices, and $\sigma(w)=\max(0,\min(1,w)),\forall w\in\{W_{M,i},W_{A,i}\}$ is utilized to ensure the weight matrices’ elements are within the range of $[0,1]$. $W_{M,i}$ and $W_{A,i}$ are initialized as all-ones matrices, and then iteratively get updated during training on local data
@@
where $\eta$ is the learning rate. During the update process of $W_{M,i}$ and $W_{A,i}$, Eqs. (10-11) and (12-13) are alternately executed for the updates of $(\hat{\phi}_{i}^{t},\hat{\bar{\phi}}_{i}^{t})$ and $(W_{M,i},W_{A,i})$. Upon convergence, the initial parameters $(\hat{\phi}_{i}^{t},\hat{\bar{\phi}}_{i}^{t})$ for the dual-decoder in each federation round are obtained, followed by the training of the local model using Eq. (1). Note that when $t>2$, only a few iterations are conducted to acquire $(W_{M,i},W_{A,i})$.